A survey looked at the value of recreational sports on college campuses. One of the questions asked each student to rate the importance of recreational sports to college satisfaction and success. Responses were on a 10-point scale with 1 indicating total lack of importance and 10 indicating very high importance. The following table summarizes these results: Class Mean score Freshman Sophomore Junior Senior 726 536 595 437 7.6 7.6 7.5 7.3 (a) To compare the mean scores across classes, what are the degrees of freedom for the ANOVA F statistic? numerator df denominator df (b) The MSG 9.815. If sp- 2.12, what is the F statistic? (Round your answer to two decimal places.)
Answer #1

a)

numerator df =categories(k)-1=4-1=3

denominator df =N-k=(726+536+595+437)-4=2290

b)

F =MSG/sp2 =9.815/2.122 =2.18